OpenStudy (nerdsarecool):

Elsa has spent $22 so far on gifts. The additional amount she will spend will be less than $15 . What are the possible total amounts she will spend? Use c for the total amount (in dollars) she will spend. Write your answer as an inequality solved for c .

OpenStudy (study312):

I am thinking 15 + c = 22 what do you think @563blackghost?

OpenStudy (nerdsarecool):

Well i will just see if that is the answer ok

OpenStudy (nerdsarecool):

No it is the wrong answer

OpenStudy (study312):

or maybe 15 + c \[\le 22\]

OpenStudy (study312):

you get a 2nd try?

OpenStudy (nerdsarecool):

I can not do that one because if i get it wrong then i will have to do another question

OpenStudy (nerdsarecool):

But if you are positive i will try just say are you positive

OpenStudy (nerdsarecool):

Are you positive or no

OpenStudy (study312):

LegendarySadist take the wheel

OpenStudy (study312):

not really but Legendary can help you

OpenStudy (anonymous):

\[\large \sf Amount~spent+c\leq Amount ~spent + Total~left~to~be~spent\]

OpenStudy (nerdsarecool):

Wait whatttttttttttttttttttttttttttttttttttttt

OpenStudy (study312):

fill in the blank of what Legendary just posted

OpenStudy (anonymous):

Just look at the problem and find which ones are equal to the values I have given you and you'll be all good

OpenStudy (nerdsarecool):

Ok thank you let me try to do it

OpenStudy (anonymous):

Nvm c is representing something else

OpenStudy (nerdsarecool):

Oh ok

OpenStudy (anonymous):

Actually I'm not sure. My original post makes the most sense, but it's worded horribly so I'm not 100% sure.

OpenStudy (anonymous):

The question is worded horribly I mean

OpenStudy (nerdsarecool):

I get the thing now thanks for helping me

OpenStudy (anonymous):

Well as long as you get it (The more I try to understand it, the less I get it tbh lol)

OpenStudy (anonymous):

What they might mean is they want you to arrange it like this \[\large \sf 22 \leq c \leq 37\] but the wording is terrible so I don't really know